Solve for x : 4x – 3x - 1/2 = 3x + 1/2  - 22x – 1

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. 4x – 3x – 1/2 = 3x + 1/2 - (22)x/2
⇒ 4x – 3x/√3 = 3x √3 – 4x/2
⇒ 3/2 4x = 3x (√3 + 1/√3) ⇒ 3/2 4x = 3x 4/√3
⇒ (4/3)x – 3/2 = 1 ⇒ x – 3/2 = 0
⇒ x = 3/2